OpenStudy (anonymous):

The coordinates of the center of a circle with equation (x − 6)2 + (y + 3)2 = 72 are _____

OpenStudy (yrelhan4):

Equation of circle with centre x1,y1 and radius r is (x − x1)^2 + (y -y2)^2 = r^2

OpenStudy (anonymous):

whats the final answer

OpenStudy (yrelhan4):

What do you think?

OpenStudy (anonymous):

-6,3

OpenStudy (yrelhan4):

as the centre is x1,y1 when the equation is (x − x1)^2 + (y -y2)^2 = r^2.. so, when the equation is (x − 6)2 + (y + 3)2 = 72 , centre would be 6,-3..

OpenStudy (yrelhan4):

@colepresc Well if you like any answer, click "best Response". :)